# RealitySummary: Exploring On-Demand Mixed Reality Text Summarization and Question Answering using Large Language Models

### v2

@@ 3.2. Results
We identified five high-level categories: summarize, compare, augment, extract, and navigate. The summarize category focused on condensing text for quick reference, ranging from broad document summaries to personalized summaries for different comprehension levels. Summaries were seen as useful reference tools, allowing users to quickly review content without losing focus. The compare category emphasized transforming textual and numerical data into visual formats like tables, graphs, and timelines, which could dynamically update based on user interactions, making comparisons more intuitive. The augment category explored ways to enhance document content with external or AI-generated data, such as adding maps and images or converting 2D figures into 3D models. Advanced ideas included visualizing the importance of academic citations or integrating external information to provide richer context. The extract category focused on enabling users to pull out and store key elements, such as keywords, figures, or quotes, for future reference. Lastly, the navigate category proposed various methods to improve document navigation, such as progress indicators, accessible tables of contents, and using AR to make documents searchable. These insights, alongside considerations of the potential and challenges of AI-driven tools, informed the development of our system.
4. Study 1 RS-Doc: RealitySummary First Prototype and Usability Study
We implemented the first version of our system on a Microsoft Hololens 2 and conducted a preliminary user evaluation. In this section, we present initial insights from using on-demand, mixed reality document summarization and discuss the associated limitations and design implications.
4.1. Implementation
Figure 4. RS-Doc Hololens 2 architecture.
Figure 5. Examples from our RS-Doc Hololens implementation.
The system design for the HoloLens version involves capturing documents using the HoloLens 2’s internal camera and extracting text with Google Cloud OCR. The extracted text is processed using GPT-4 via OpenAI’s API to identify key information, while spaCy’s named entity recognition (NER) module assists in identifying important entities for further web searches. The system employs Vuforia for markerless image tracking and the Mixed Reality Toolkit 3 (MRTK3) SDK for content rendering. Users interact with the system through voice commands processed via Google’s speech-to-text API, allowing them to modify or query visualizations. The system captures images periodically for OCR, with an option for manual activation, and tracks documents by generating dynamic image targets, ensuring accurate document tracking and augmented content rendering.
Informed by the elicitation study, the system supports several document enhancements aimed at improving the reading experience. These enhancements include text summarization, achieved by sending the current page’s text to GPT-4 for a concise summary, limited to 150 words for easier digestion. Comparison tables are generated on demand based on the document parsing, allowing for quick comparisons within the document. Timelines of events are created by prompting GPT-4 to generate a sequence of summarized events. A keyword list is extracted using GPT-4 and spaCy, with spaCy handling named entities for web searches and keyword highlighting. Summary highlighting allows users to obtain brief summaries or additional information on specific phrases, sentences, or paragraphs, which are then visually highlighted. Information cards provide further details on keywords, people, places, or images, leveraging Google Search, Image APIs, and Mapbox for generating maps. All these features are integrated into the HoloLens environment, offering real-time, interactive content within the mixed reality experience.
4.2. Method
The usability study involved 12 participants (P1-P12, 6 male and 6 female, aged 19-35) recruited via word-of-mouth and compensated \$15 for their one-hour participation. The study aimed to evaluate the system’s usability, explore participants’ preferences for different document enhancement strategies, and identify limitations for future iterations with real-world use in mind. Conducted in a lab environment with Microsoft HoloLens 2, participants were first introduced to AR and the HoloLens 2, followed by an overview of the system and its document enhancements. They completed three reading tasks using different document formats (physical, digital, and a book) and utilized the system’s enhancements as they preferred. The participants’ digital documents included books, blog posts, course assignments, Wikipedia articles, magazine articles, and news articles viewed on various devices.
Throughout the study, participants were encouraged to think aloud, and their experiences were documented. Following the tasks, semi-structured interviews and a Likert scale questionnaire, including a System Usability Scale (SUS), were conducted. Thematic analysis was performed on the transcriptions of the think-aloud sessions and interviews, resulting in the identification and refinement of key themes.
4.3. Results
@@
Overall, participants responded positively to RealitySummary, with most participants agreeing that the prototype was easy to use (M=4.3/5, SD=0.98), easy to understand (M=4.5/5, SD=0.67), and well-integrated with the task (M=4.3/5, SD=0.77) with 5-point Likert-scale questionnaire. Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with comprehension (M=4.6/5, SD=0.514), were relevant to the document content (M=4.7/5, SD=0.65), improved comprehension speed (M=4.4/5, SD=0.79), and augmented their reading experience (M=4.7/5, SD=0.65). Finally, the System Usability Scale (SUS) returned an overall composite score of 71, suggesting a reasonable level of usability for an early-stage prototype.
4.4. Insights
@@
Our analysis identified several themes related to participants’ experience in comprehending and navigating documents. We also explore participants’ comparisons of the visual and text-based summarization techniques as well as preferences for various augmentation techniques.
Enable Multi-Angle Content Understanding for the Current Page
RealitySummary’s document enhancements can significantly improve readers’ comprehension. Several participants (P7, P8, P11, P12) highlighted that features such as summaries, timelines, keywords, images, and people cards provided a more comprehensive understanding of the document, allowing them to engage with the content from multiple angles and perspectives. All participants mentioned that RealitySummary enabled them to grasp content more quickly than with their usual methods, such as visually scanning a physical book or manually searching for timelines on a phone or PC. For instance, after reviewing the first magazine, P12 remarked, I haven’t read this magazine thoroughly, but it seems these features can significantly help me understand what it’s about in less time.” Interestingly, several participants appreciated that the summaries and keywords were drawn only from the currently visible page rather than from the entire document. P5 reported, I like the page-by-page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document.” Similarly, P6 emphasized, “I like that it (the system) gives you keywords for each page and not for the whole document.
Facilitate Efficient, Top-Down Navigation
@@
Both structured overviews and in-context highlights provide useful entry points that make it easier to navigate documents from a top-down perspective, even within individual pages. Broadly, participants spoke positively about how RealitySummary helped them navigate documents. Four participants (P3, P4, P6, P8) specifically mentioned using structured navigation elements like timelines to better understand the structure and flow of the text, with P4 noting that the timeline helped me organize and navigate my thoughts very quickly while reading the text.” In fact, ten participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining the high-level ideas and terms, then transitioning into reading the lower-level content as needed. P8 highlighted, “[the enhancements] give you a foundation before you actually jump inside [the document].” This approach enabled participants to prioritize their reading, allowing them to “focus on things that you’re more unsure about” (P1).
Support Quick Understanding with Visual Aids
@@
Both visual and text-based enhancements offer distinct benefits to readers—visual aids provide quick, high-level overviews, while text summaries capture more complex details. Some participants preferred text-based features for gaining a deeper understanding of specific topics, while others found that visual elements improved their comprehension speed, engagement, and retention. All 12 participants rated the visual summarization techniques as helpful, with eight specifically highlighting the usefulness of images and bio cards. Several participants (P1, P4, P6, P7, P9, P10, P11) noted that visual augmentations allowed them to grasp the document’s content more quickly compared to reading text alone, with P11 stating, “Seeing the visuals helped me grasp what’s happening in the document right away.” Participants (P7, P8) also emphasized that visuals enhanced their engagement with the material, with P7 commenting, I’m not a big fan of reading, so having visuals helps me engage with the content.” Additionally, participants highlighted the potential of visuals to improve memorability, with P8 noting that for historical documents, people cards helped put a face to a name.” Similarly, P7 observed that while text-based summaries aided understanding, “graphical elements help me remember it for a longer time.”
4.5. Limitations
Slow and Unreliable Document Tracking
Document tracking enables unique features like keyword highlighting and around-document placement. However, the primary limitation of the current prototype is the instability of its document tracking. Overall, participants reported dissatisfaction with the system’s slow, inaccurate, and unreliable tracking capabilities. Unlike the OCR process, document tracking involves a heavier and slower workflow, requiring rectangle detection, image extraction, and image upload for Vuforia processing. Even with a perfect image target, tracking performance varies based on several factors, such as lighting conditions, hand occlusion, and viewing angles. Furthermore, while the tracking performed better for visually rich documents, it struggled with text-heavy documents, as it does not have any distance features, making consistent tracking difficult over time. Even when tracking functioned properly, participants frequently experienced disruptive flickering, which negatively affected the user experience. Additionally, the fixed placement of augmented information on documents posed a challenge for long-term use, as participants expressed the need for more flexibility in adjusting the location of the augmentations.
@@ 4.5. Limitations
Fixed Prompt and Augmentation Types
Another key issue was the inability of participants to query the generated summaries. The fixed prompt approach used for generating the different types of cards limited the system’s flexibility in addressing broader use cases and contexts. Several participants (P1, P3, P5) expressed frustration with this constraint, as they had specific questions they wanted to ask but were restricted to the preset summary types provided. While these different types of augmentations offered different benefits, those advantages were only realized when the generated content was relevant and accurate. Participants also noted that specific aids, like timelines, comparison tables, and keyword lists, were not always helpful or contextually appropriate. On the other hand, participants generally appreciated the broader usefulness of the general summaries and visual information cards, which were more adaptable to varying contexts.
Applicability Beyond Controlled Environments
Finally, the controlled nature of the study limits its ability to assess the broader applicability and real-world uses. While the initial findings are encouraging, they do not fully reflect how the system might perform in diverse, unstructured environments. As a result, transitioning to in-the-wild studies will be essential to understanding how document augmentations work in everyday scenarios and to gain insights from prolonged, regular use in real-world settings.
In conclusion, a more refined version of the system, featuring flexible augmentation placement and query-able summaries, is essential for better evaluating its effectiveness in real-world scenarios.
5. Study 2 RS-WILD: In-the-Wild Study with the Second Prototype
Figure 6. Examples from our in-the-wild study.
To explore real-world use cases, we conducted an in-the-wild study with eleven participants. Unlike the previous usability study, which focused on standardized reading tasks in a controlled lab setting, this study aimed to examine the versatility and applicability of our system in uncontrolled environments and with diverse reading materials. The goal was to uncover unique advantages and potential use cases for using mixed reality and AI-generated assistants in everyday settings. For this study, the system was deployed on an Apple Vision Pro, offering considerably higher rendering quality and a broader field of view compared to the Hololens 2. This contributes to better readability, particularly for text summarization, by using a less distractive transparent background. Based on the user feedback from the usability study, the user interface was simplified with a single text box and an image slot, which could accommodate any response type given the user’s custom query. This feature would better suit diverse applications and minimize the intrusiveness of additional content in daily use.
@@ 5.1. Implementation
The system design for the Vision Pro follows a similar approach to the Hololens but with key adjustments. Since the Vision Pro does not yet allow direct access to the camera view, we implemented a workaround for text extraction. First, we screencast the Vision Pro to an external iPad, which is then connected to a MacBook Air to stream the feed via QuickTime Player’s mirroring feature. A web-based interface on the Mac captures the live stream from the Vision Pro. Using the captured feed, we employed Google OCR to extract text, which was then processed by GPT-4 via the OpenAI API. The Vision Pro application, developed in Swift, featured a single movable card with an image/map slot. After querying the LLM, data was transmitted from the web-based interfaces to Vision Pro via Firebase’s real-time database.
Additionally, a separate web interface with speech recognition was run on the user’s smartphone, allowing them to query the system via voice (using Web Speech Recognition) and ask follow-up questions. This enabled users to request specific summary types or submit open-ended queries about the text they were reading. The web-based interfaces, built with React.js and hosted on Glitch, were accessed by participants through their phones.
5.2. Method
We recruited 11 participants (W1-W11, 7 male and 4 female; ages 19-32 years) from the local community. They were compensated \$25 for participating in a 1.5-hour study, which included a one-hour experiment and a 30-minute interview. The study took place in various locations at a local university, such as a library, cafeteria, store, hallway, classroom, and multiple labs.
@@ 5.2. Method
Before the experiment, we collected participants’ demographic information and reading habits, which were input into the system as prior knowledge for ChatGPT. Participants familiarized themselves with the Apple Vision Pro by completing the eye-tracking calibration setup. Unlike structured tasks or prepared reading materials, participants were encouraged to use the system with any text they encountered during the study, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment. Throughout the one-hour experiment, participants were encouraged to think aloud about their experiences. After the experiment, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the one used in the earlier usability study. Both the interview and questionnaire took approximately 30 minutes. As in the previous usability study, we analyzed participants’ questionnaire responses and reflections to identify high-level trends and opportunities for augmented reading tools.
@@ 5.3. Results
We gathered feedback from participants on various aspects, including the reading experience (immersiveness, understanding, distraction), perceptions of the system (content and query comprehension, trustworthiness, and privacy concerns), and overall usability (content placement and comfort using the MR headset). In general, participants reported feeling more immersed and interested while reading (M=3.5/5, SD=1.04), and they perceived the system as effectively interpreting their queries (M=3.3/5, SD=1.00). Besides, their responses on the content placement within the system were also overall positive (M=3.4/5, SD=1.29).
Figure 7. In-the-wild study questionnaire responses.
Participants had varied opinions on whether the system became distracting throughout the reading process (M=2.6/5, SD=1.37) and whether they tended to trust the accuracy and reliability of contents provided by the system (M=2.8/5, SD=1.25). Privacy concerns were also noted by the majority (6/11), and discomfort using the headset was prevalent among the respondents (8/11). These findings are detailed in Figure 7.
5.4. Insights
@@
Types of Reading Content
@@
The in-the-wild study demonstrated our system’s broad utility, revealing diverse exploratory use cases across various reading materials. Participants engaged with a variety of books, spanning textbooks, encyclopedias, fiction, and non-fiction. The system was also employed for different types of documents, including academic papers, handouts, and class assignments, as well as digital content accessed through smartphones and laptops, like social media posts, online news articles, and e-commerce websites. Beyond conventional reading materials, participants also used our system with everyday objects and texts in the environment, such as restaurant menus, printer instructions, product nutrition labels, furniture assembly guides, posters, and signage on trash bins and advertisements. Notably, some participants even applied our system to their handwritten notes from lectures.
Benefits of Always-On AI Assistant
The exploration of such diverse reading materials can be attributed to the always-on capability of the mixed reality reading assistant, as opposed to screen-based interfaces. A key benefit of this feature is its support for implicit inputs, which dramatically lowers the barrier to asking questions of LLMs. This automatic, continuous content capture broadens the range of applications beyond what participants initially envisioned. For instance, W4 mentioned they had never thought to type a question into ChatGPT when looking at a product label. Even for common reading tasks, they appreciated that the summary was automatically generated and updated in the background as they flipped through pages in a library. Overall, the always-on, implicit assistant made it effortless to integrate AI into daily tasks, expanding potential use cases and drastically reducing the explicit effort required from users.
Unique Affordances of Spatial and Tangible Interactions
@@
Participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration. In a library setting, for example, participants could naturally navigate through bookshelves (spatial exploration), physically pick up a book (tangible interaction), and browse through its pages to automatically receive summaries (always-on assistance). W5 and W10 even leveraged our system to provide summaries or recommendations for entire collections of books on a bookshelf, not just content from a single page. Furthermore, W1 demonstrated the capability to synthesize information across multiple books on a table, requesting summaries or connections between various books and pages. Such observations highlight how mixed reality interfaces enable spatial and tangible interactions that go beyond the limitations of traditional screen-based interfaces, which can only capture and summarize visible content.
Unique Ways of Using AI-Generated Content
@@
We observed several unique and unexpected use cases that took advantage of the LLM’s capacity for free-form questioning and answering. For instance, when reading a physics textbook, W8 asked the system to generate ten quizzes to help him understand quantum mechanics, transforming the textbook into a more active and personalized learning tool. Other common usages included translations (W4, W5, W6), learning aids (W4, W8), and personalized content recommendations (W5, W10). For example, W8 applied the system to his handwritten calculus notes to find answers to particular questions. Beyond traditional reading applications, the AI’s capabilities supported practical decision-making; W4 compared nutritional information of products on store shelves and menu items to determine healthier options, while W5 used it for guidance on waste disposal and printer operations. The LLM’s ability to process a broad spectrum of inquiries clearly expanded the system’s applicability and utility.
Proactive Summaries vs On-Demand Question Answering
@@
Our system offers both 1) proactive assistance through automatic summary generation and 2) on-demand assistance where users manually ask questions. Participants shared their insights on these two functionalities by comparing the benefits and limitations. Proactive summaries were appreciated for providing a useful initial overview of unfamiliar topics. Preferences for proactive summaries varied with reading habits. For example, infrequent readers like W6 and W10 found these proactive summaries helpful in deciding whether to read the content or understand it without reading the full text. On the other hand, W10, who is a more frequent reader, found the constant summary updates unnecessary. In contrast, on-demand question answering allowed for a more focused engagement on specific topics. For instance, W1 enjoyed the “rapid-fire” questioning with a physics textbook, asking more questions than with traditional ChatGPT saying ”You have to describe it properly you have to specifically say what you’re asking about what you’re reading, so that kind of all of this abstracted in this system is like frees up your mind to just focus on the curiosity part Like rapid-fire questions and get rapid-fire answers”. W2 valued how questioning was woven into reading, preserving thought flow without the disruption of toggling between reading and ChatGPT interfaces. Overall, participants favored the on-demand feature due to its applicability to diverse situations, yet many appreciated having both options for different purposes: proactive features for implicit skimming and on-demand for addressing more specific questions.
Potential Risks of AI-Driven Augmented Reading
Participants also expressed several concerns about potential risks, notably privacy issues due to the possibility of constant surveillance. Despite this concern, most participants were willing to continue using the system, if they could control when it was active. For example, W1 expressed the desire to capture all text throughout her entire life, allowing for an extensive searchable archive of everything she has encountered. Additionally, participants also mentioned the risk of decreased motivation and skill development. For example, W6 mentioned that relying too heavily on the system for translations could reduce the motivation to learn new languages. Similarly, W11 worried that immediate access to answers for every question might contribute to a broader intellectual complacency. Overall, participants wanted the flexibility to choose when and how the system would be used to mitigate these risks.
Trust of AI-Generated Content
Many participants approached the system’s responses with caution, often verifying the answers with the reading material. They highlighted the necessity of consulting additional sources for in-depth or technical readings when depending on the system. There was a general agreement among participants about their reluctance to fully trust an always-on reading assistant in their daily activities. Conversely, we noted some intriguing discussions regarding how participants perceive our system, either as a mere tool or more like a personal companion. While most participants engaged with the system as a conventional chatbot, some regarded it more as a personal companion than a mere tool. Notably, W3 and W5 described the system as a friend for discussing and inquiring about their readings. The system’s ability to provide always-on, immediate responses facilitates this personification.
5.5. Limitations
Overall, the system design of the second prototype enhanced its applicability and stability for real-world use. Compared to the first prototype on the Hololens 2, the Vision Pro’s superior rendering quality and broader field of view provided a better user experience. Additionally, the simple system design greatly improved usability despite the absence of document tracking. While the second prototype showed great potential, we also identified several feature-level limitations, along with inherent study-related constraints.
Lack of History and Recall Support
While the minimalist design effectively reduces visual clutter and conserves screen space, it also limits the system’s overall utility. Although users appreciated the clean and simple interface, many expressed frustration with the lack of options to save or revisit previously generated responses. Several participants suggested that a history feature or the ability to store responses for future reference would substantially enhance the user experience.
Challenges with Screen Viewing for Extended Use
Many participants raised concerns about the difficulty of clearly viewing phone and computer screens through the Vision Pro without dramatically enlarging the text. While this worked for short-term tasks like quickly reviewing an assignment, participants found it frustrating for longer viewing periods, particularly when using laptops. Since much of their work is done on computer screens, this limitation poses a challenge for assessing the system’s effectiveness in supporting realistic, work-related tasks.
Need for a Longer-Term Study
Our in-the-wild study was inherently limited by its short-term, public use — with each session lasting only one hour per participant. As a result, participants did not have the opportunity to fully engage with the system in their personal workflows. For instance, due to privacy concerns related to screen recording, participants rarely used our system for more personalized content, such as emails, messages, or calendars. Furthermore, since participants were encouraged to explore various scenarios, they tended to experiment broadly rather than focusing on any specific task in depth. While the in-the-wild study provided valuable insights into real-world use cases, it only captured short-term interactions in public spaces. There is still much to learn about the sustained, regular use of reading assistants in daily life, particularly in how users interact with information over time, revisit previous queries, or support ongoing tasks.
6. Study 3 RS-Diary: Diary Study with the Third Prototype
To explore the long-term use cases, we developed a third iteration, incorporating feedback from the in-the-wild study. This version introduced a multiple responses feature and added the ability to view or mirror the content on an external device, addressing the readability issues for the long-term use of the Vision Pro.

### v3

@@ 3.3. Results
Overall, participants responded positively to RealitySummary, with most participants agreeing that the prototype was easy to use (M=4.3/5, SD=0.98), easy to understand (M=4.5/5, SD=0.67), and well-integrated with the task (M=4.3/5, SD=0.77) on a 5-point Likert-scale questionnaire. Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with perceived comprehension (M=4.6/5, SD=0.514), were relevant to the document content (M=4.7/5, SD=0.65), perceived improvements in comprehension speed of the document (M=4.4/5, SD=0.79), and augmented their reading experience (M=4.7/5, SD=0.65). Finally, the System Usability Scale (SUS) returned an overall composite score of 71, suggesting a reasonable level of usability for an early-stage prototype.
3.4. RS-Doc Observations
@@
Our analysis identified several themes related to participants’ experience comprehending and navigating documents. We also explore participants’ comparisons of the visual and text-based summarization techniques as well as their preferences for various augmentation techniques.
Page-Level Summaries and Keyword Lists Support Quick Comprehension.
Since our RS-Doc system can only capture the currently visible page, our summaries are inherently focused on page-level content rather than the entire book. Interestingly, when asked about their experiences, several participants expressed appreciation for this feature. P5 reported, “I like the page-by-page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document.” Similarly, P6 emphasized, “I like that it (the system) gives you keywords for each page and not for the whole document.
Structured Overviews can Facilitate Efficient Document Navigation.
@@
Both structured overviews and in-context highlights provide useful entry points that make it easy to navigate documents from a top-down perspective, even within individual pages. Four participants (P3, P4, P6, P8) specifically mentioned using structured navigation elements like timelines to better understand the structure and flow of the text, with P4 noting that “the timeline helped me organize and navigate my thoughts very quickly while reading the text.” In fact, ten participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining the high-level ideas and terms, then transitioning into reading the lower-level content as needed. P8 highlighted, “[the enhancements] give you a foundation before you actually jump inside [the document].” This approach enabled participants to prioritize their reading, allowing them to “focus on things that you’re more unsure about” (P1).
Visual aids improve engagement and memorability.
@@
Both visual and text-based enhancements offer distinct benefits to readers—visual aids provide quick, high-level overviews, while text summaries capture more complex details. Some participants preferred text-based features for gaining a deeper understanding of specific topics, while others found that visual elements improved their comprehension speed, engagement, and retention. All 12 participants rated the visual summarization techniques as helpful, with eight specifically highlighting the usefulness of images and bio cards. Several participants (P1, P4, P6, P7, P9, P10, P11) noted that visual augmentations allowed them to grasp the document’s content more quickly compared to reading text alone, with P11 stating, “Seeing the visuals helped me grasp what’s happening in the document right away.” Participants (P7, P8) also emphasized that visuals enhanced their engagement with the material, with P7 commenting, I’m not a big fan of reading, so having visuals helps me engage with the content.” Additionally, participants highlighted the potential of visuals to improve memorability, with P8 noting that for historical documents, people cards helped “put a face to a name.” Similarly, P7 observed that while text-based summaries aided understanding, “graphical elements help me remember it for a longer time.”
3.5. Opportunities for the Second Iteration
Building on these observations, we identified a number of opportunities to explore in our second iteration.
Reader-Controllable Placement Rather Than Document Tracking
An intriguing finding that might contradict previous work is the role of document tracking. For on-demand document enhancement, where dedicated markers or image targets cannot be prepared in advance, unreliable or slightly jittery document tracking often leads to frustration. For example, while HoloDoc (Li et al., 2019) and Dually Noted (Qian et al., 2022) relied on dedicated markers or pre-defined image targets, enabling much more reliable document tracking, our workflow cannot rely on these methods. Instead, we need to rely on complex and unreliable image extraction workflows. Moreover, document tracking performed well only with visually rich documents. Participants expressed frustration when dealing with text-heavy documents, particularly under challenging conditions such as poor lighting, hand occlusion, or extreme viewing angles. Participants also expressed the need for flexibility in the placement. Therefore, while we acknowledge the benefits of document tracking, it is worth exploring reader-controlled placement as a less frustrating alternative.
@@ 3.5. Opportunities for the Second Iteration
Querying and Contextual Flexibility
Participants highlighted the potential for enhancing the system by enabling queries on the generated summaries, which could expand its adaptability to broader use cases and contexts. The current fixed prompt approach for generating different types of cards suffices for general reading but could be further refined to address more specific reader questions. Several participants (P1, P3, P5) envisioned a more conversational system where they could ask targeted questions beyond the preset summary types. While participants found general summaries and visual information cards broadly useful and adaptable to varying contexts, they also identified opportunities to enhance aids with specific queries for timelines, comparison tables, and keyword lists that better align with their content.
Overall, the controlled nature of the study limited its ability to assess broader applicability and real-world uses. While the initial findings are encouraging, they do not fully reflect how AR summarization and question-answering tools might perform in diverse, unstructured environments. As a result, transitioning to in-the-wild studies is essential to understanding how these augmentations might work in everyday scenarios.
4. RS-Wild: In-the-Wild Study with the Second Prototype
To explore these kinds of real-world use cases, we developed a second prototype (RS-Wild) that allowed us to study summarization and question-answering in-the-wild. To address performance issues with the HoloLens 2 and support more free-ranging use, we instead implemented our RS-Wild prototype for the Apple Vision Pro. This provided higher rendering quality for cards, as well as a broader field of view. Building on observations from our prior study, we also simplified the interface—replacing the suite of document-anchored cards from RS-Doc with a single reader-positionable card that readers could dynamically update using spoken queries.
@@ 4.1. Implementation
Our RS-Wild implementation for the Vision Pro follows a similar approach to the HoloLens but with a few key adjustments. Since the Vision Pro does not yet allow direct access to the camera feed, we implemented a workaround for text extraction. First, we screencast the Vision Pro to an external iPad, which is then connected to a MacBook Air to stream the feed via QuickTime Player’s mirroring feature. A web-based interface on the Mac captures the live stream from this mirror and by proxy, from the Vision Pro. Using the captured feed, we employed Google OCR to extract text, which was then processed by GPT-4 via the OpenAI API. After querying the LLM, data was transmitted from the web-based interfaces to the Vision Pro via Firebase’s real-time database. Voice commands were handled via a separate interface (using the Web Speech API and React.js and hosted as a progressive web app) loaded on the reader’s smartphone.
Interactions
RS-Wild allows readers to request summaries and pose free-form questions about any text visible in their environment by clicking on a mic button on the smartphone interface and asking a question or clicking a “summaries” button. This then triggers the OCR and LLM pipeline. Responses to the most recent request are displayed on a single reader-positioned card which can be placed anywhere in the environment. This added flexibility allowed viewers to recreate any of the individual cards available in RS-Doc as well as pose new unanticipated questions and request new summary types. Using a single manually-positioned card with a transparent background also allowed readers to position that response or summary in more useful locations in their environment—placing it both close to and apart from documents as appropriate for their current task. Moreover, the user can choose to keep a history of all the text that has been captured in their session, allowing them to pose questions with a little context. Additionally, the card follows the user’s position in MR space, so they always have access to it even while walking or changing locations.
4.2. Study Method
To explore question-answering and summarization in-the-wild we recruited a new cohort of 11 participants (W1-W11, 7 male and 4 female; ages 19-32 years) from our university through snowball sampling and university advertisements. All participants were university students accustomed to reading in academic environments, which met our inclusion criterion of being able to read documents of varied comprehension levels. No additional exclusion criteria were applied. We compensated participants \$25 CAD for the 1.5-hour study, which included a one-hour experiment and a 30-minute interview. The participants were first called to our lab, briefed on the procedure, and given the opportunity to pick one or multiple places around the university campus (such as a library, cafeteria, store, hallway, classroom, or various labs) to explore with the system. They were allowed to freely explore any place in the campus with the headset and the study conductor would follow them.
@@ 4.2. Study Method
Before the experiment, we collected participants’ demographic information and reading habits, which we then input into the system as prior knowledge for GPT-4. Participants familiarized themselves with the Apple Vision Pro by completing the eye-tracking calibration setup. Instead of structured tasks or prepared reading materials, we encouraged participants to use the system with any text they encountered during the study, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment. We also directed participants to explore various locations at a local university, rather than using spaces prearranged by the experimenters. Throughout the one-hour experiment, an experimenter accompanied the participants, observing their usage and encouraging them to think aloud about their experiences. After the experiment, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the one used in Study 1.
As in the previous usability study, we analyzed participants’ questionnaire responses and transcripts to identify high-level trends and opportunities for augmented reading tools. We performed a reflexive thematic analysis (Clarke and Braun, 2014) using transcripts of both the think-aloud sessions and interviews, from which we identified and refined key themes.
@@ 4.3. Results
We gathered feedback from participants on various aspects, including the reading experience (immersiveness, understanding, distraction), perceptions of the system (content and query comprehension, trustworthiness, and privacy concerns), and overall usability (content placement and comfort using the MR headset). In general, participants reported feeling more immersed and interested while reading (M=3.5/5, SD=1.04), and they perceived the system as effectively interpreting their queries (M=3.3/5, SD=1.00). Their responses on the content placement within the system were also positive overall (M=3.4/5, SD=1.29).
Participants had varied opinions on whether the system became distracting throughout the reading process (M=2.6/5, SD=1.37) and whether they tended to trust the accuracy and reliability of contents provided by the system (M=2.8/5, SD=1.25). Privacy concerns were also noted by the majority (6/11), and discomfort using the Vision Pro headset was prevalent among the respondents (8/11).
4.4. RS-Wild Observations
@@
Diverse Reading Content
@@
Participants engaged with a variety of books (spanning textbooks, encyclopedias, fiction, and non-fiction) as well as other types of documents (including academic papers, handouts, and class assignments) and diverse digital content accessed through smartphones and laptops (like social media posts, online news articles, and e-commerce websites). Beyond conventional reading materials, participants also used our system with everyday objects and text in the environment, such as restaurant menus, printer instructions, product nutrition labels, furniture assembly guides, posters, signage on trash bins, and advertisements. Notably, some participants even applied our system to their handwritten notes from lectures.
Incorporating Spatial and Tangible Interactions
@@
Participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration. In a library setting, for example, participants could naturally navigate through bookshelves (spatial exploration), physically pick up a book (tangible interaction), and browse through its pages to automatically receive summaries. W5 and W10 even leveraged our system to provide summaries for entire collections of books on a bookshelf, not just content from a single page. Furthermore, W1 demonstrated the ability to synthesize information across multiple books on a table, requesting summaries or connections between various books and pages. Such observations highlight how mixed reality interfaces enable spatial and tangible interactions that go beyond the limitations of traditional screen-based interfaces, which can only capture and summarize visible content.
Creative Uses of AI-Generated Responses
@@
We observed several unique and unexpected use cases that took advantage of the LLM’s capacity for free-form questioning and answering. For instance, when reading a physics textbook, W8 asked the system to generate ten quizzes to help him understand quantum mechanics, transforming the textbook into a more active and personalized learning tool. Other common uses included translation (W4, W5, W6), creating learning aids (W4, W8), and generating personalized content recommendations (W5, W10). For example, W8 applied the system to his handwritten calculus notes to find answers to particular questions. Beyond traditional reading applications, the AI’s capabilities supported practical decision-making. For example, W4 compared nutritional information of products on store shelves and menu items to determine healthier options, while W5 used it for guidance on waste disposal and printer operations. The LLM’s ability to process a broad spectrum of inquiries clearly expanded the system’s applicability and utility.
Proactive Summaries vs Reader-Driven Question Answering
@@
Our system offers both 1) proactive assistance through automatic summary generation and 2) reader-driven assistance where readers manually ask questions. Participants shared their insights on these two functionalities by comparing their benefits and limitations. Readers appreciated how proactive summaries could often provide a useful initial overview of unfamiliar content. Preferences for proactive summaries varied with reading habits. For example, infrequent readers like W6 and W10 found these proactive summaries helpful in deciding whether to read the content at all. On the other hand, W10, who is a more frequent reader, found the constant summary updates unnecessary. In contrast, reader-driven question answering allowed for a more focused engagement on specific topics. For instance, W1 enjoyed “rapid-fire” questioning with a physics textbook, noting that it encouraged them to ask more questions than they would with traditional ChatGPT. W2 valued how questioning was woven into reading, preserving thought flow without the disruption of toggling between reading and chat or search interfaces. Overall, participants favored the reader-driven mode due to its applicability to diverse situations, yet many appreciated having both options for different purposes—proactive features for implicit skimming and reader-driven for addressing more specific questions.